I thought I was getting the hang of this but definitely not sure now-
x^2-16x+4=0. Please show your work PLEASE! I need the value of the discriminate, give the number and the number of roots/solutions/zeros. I thought that I was to ad x^2+16+4 but I am so confused that I don't even know any more.
Again please show your work so that I may understand the processes involved.

Answer by Edwin McCravy(20055)

The discriminant is the part of the quadratic formula under the
square root, b²-4ac.  

If it's positive, there are two real solutions
If it's zero, there is 1 real solution
If it's negative, there are no real solutions 
(but there are two imaginary solutions)

x² - 16x + 4 = 0

Always first compare an equation like this to this:

ax² + bx + c = 0

There is an understood 1 before th x², so think of your equation as

1x² - 16x + 4 = 0

Then a = 1, b = -16 and c = 4

The formula for the discriminant is 

D = b² - 4ac

Substitute in the discriminant formula:

D = (-16)² - 4(1)(4)

D = 256 - 16

D = 240.  That's positive so there are two real solutions.

Edwin